### v1

# Shortest Paths in a Weighted Simplicial Complex

Abstract
Simplicial complexes are extensively studied in the field of algebraic topology. They have gained attention in recent time due to their applications in fields like theoretical distributed computing and simplicial neural networks. Graphs are mono-dimensional simplicial complex. Graph theory has application in topics like theoretical computer science, operations research, bioinformatics and social sciences. This makes it natural to try to adapt graph-theoretic results for simplicial complexes, which can model more intricate and detailed structures appearing in real-world systems. In this article, we define the concept of weighted simplicial complex and $d$-path in a simplicial complex. Both these concepts have the potential to have numerous real-life applications. Next, we provide a novel algorithm to find the shortest paths in a weighted simplicial complex. The core principles of our algorithm align with those of Dijkstra’s algorithm for graphs. Hence, this work lays another brick for the sake of integrating graph-theoretic concepts with abstract simplicial complexes.
Keywords:
Weighted Simplicial Complex, Shortest Path, Greedy Algorithm, Dijkstra’s Algorithm
1 Introduction
Simplicial Complexes are objects that are studied extensively in algebraic topology using tools like homology and cohomology. Simplicial complexes can capture intricate topological properties, such as holes and voids, across multiple dimensions. Simplicial complexes are widely applied in algebraic topology, computational geometry, and data analysis, particularly for capturing the shape of data (e.g., in persistent homology) and for the study of spaces with more complex topological structures. On the other hand, an abstract simplicial complex is the combinatorial counterpart of a simplicial complex. Abstract simplicial complex is often studied algebraically, thus forming a relation between combinatorics and commutative algebra. We formally define abstract simplicial complexes in the next section.
Now a graph can be seen as a boiled down version of a simplicial complex. While a graph is typically a 1-dimensional object, an abstract simplicial complex can exist in any finite dimension. One of the main differences between a graph and a simplicial complex is that in a simplicial complex with $d\geq 2$, more than two $((d-1))$-dimensional simplices can be connected by a single $d$-dimensional simplex, whereas in a graph, an edge can connect at most two vertices.
Graphs are commonly used in computer science, network theory, social science, and many other fields to represent relationships between entities, such as in social networks, computer networks, and transportation systems. A graph is usually analyzed in terms of its connected components, paths, and cycles.
On the other hand, Simplicial complexes have also been used in various fields as detailed in [6, 5, 7]. For example, in [6] and [5], the results on simplicial complexes were used to prove new results in theoretical distributed computing. Distributed computing is a field of computer science that studies distributed systems. In [7], Kozlov studied the chromatic subdivision of a simplicial complex. Chromatic subdivision plays a central role in the analysis of the solvability of tasks in different computational models in computer science (see [2] for details). Simplicial neural networks (SNN) have emerged as a new direction in graph learning, extending the concept of convolutional architectures from node space to simplicial complexes, as detailed in [4, 3, 12]. The reconstruction of random simplices (see [1, 9, 8]) has emerged as a recent and significant trend in the study of random structures, gaining attention for its theoretical importance and potential applications.
Hence, to better equip abstract simplicial complexes to counter more real life problems, extending results from graphs to abstract simplicial complexes is a natural progression, as the latter provide a more versatile and detailed framework for modeling the complex structures observed in many real-world phenomena. Consequently, deeper insights into these complex structures and their underlying properties can be gained by extending graph-theoretic methods to simplicial complexes.
1.1 An Example of Real-life Application of Simplicial Complexes
Consider a town with housing complexes $A$, $B$, and $C$, where each complex contains two housing flats: $A1,A2,B1,B2,C1,C2$. A delivery boy is tasked with delivering parcels to each flat. The delivery boy can travel from the delivery office (denoted $O$) to a specific complex and deliver parcels to multiple flats within that complex.
In this setup:
  • Each flat ($A1,A2,B1,B2,C1,C2$) and the delivery office ($O$) is represented as a vertex.
  • A single edge from $O$ to a complex $A$ represents the delivery to both $A1$ and $A2$.
This situation is better modelled using a simplicial complex rather than a graph, as a simplicial complex allows for the representation of higher-dimensional relationships. Specifically, a simplicial complex can represent subsets of vertices (e.g., $\{O,A1,A2\}$) as simplices, naturally capturing the interaction of multiple entities (flats) within the same complex.
1.2 Our Contributions
The concept of given weights associated to edges in a graph has immense real-life application. With this thought, in this article, we define the concepts of a weighted simplicial complex and a $d$-path in the complex. The previously described real-life example (see Section 1.1) showcases the importance of our constructions in real-life. How to find the shortest path in a weighted simplicial complex is the question that intuitively follows. We provide a novel algorithm to find the shortest path in a weighted simplicial complex. Our algorithm is based on the same underlying principles as Dijkstra’s algorithm for graphs. We also provide correctness arguments for our algorithm along with thorough time and memory complexity analysis.
1.3 Outline
In Section 2, we provide all the notations and definitions necessary for explaining our algorithm. This section also introduces the concepts of weighted simplicial complex and $d$-path. The section ends with a description of the Dijkstra’s algorithm for graphs. In Section 3 we describe our novel algorithm to find the shortest path in a simplicial complex. The section ends with a detailed analysis of the algorithm. We conclude our study with some remarks and possible future research directions in Section 4.
2 Preliminaries
In this section, the notations and definitions that will form the building blocks for our algorithm will be discussed. This section also covers all the newly introduced concepts in this paper. Furthermore, for didactic reasons, we also explain Dijkstra’s algorithm for graphs at the end of this section. This will help the reader to relate our set-up with that of the graphs and see how our algorithm can be thought of as a higher-dimensional generalisation of Dijkstra’s algorithm for graphs.
2.1 The Set-up
We start by formally defining a finite abstract simplicial complex.
Definition 1 (Finite Abstract Simplicial Complex)
Let $\mathcal{X}_{0}$ be a finite set. A finite abstract simplicial complex $X$ on $\mathcal{X}_{0}$ is a collection of nonempty subsets of $\mathcal{X}_{0}$ such that:
  1. 1.
    For every element $x\in\mathcal{X}_{0}$, the singleton $\{x\}$ belongs to $X$,
  2. 2.
    If $\sigma\in X$ and $\tau\subset\sigma$, then $\tau\in X$.
For example, $X=\{\{1\},\{2\},\{3\},\{1,2\},\{2,3\},\{1,3\},\{1,2,3\}\}$ is an abstract simplicial complex on $\mathcal{X}_{0}=\{1,2,3\}$ (see Figure 1).
$\{1,2,3\}$
Figure 1: Example of a finite abstract simplicial complex
For any finite set $A$, let $|A|$ denote the number of elements in $A$. A set $\sigma\in X$ with $|\sigma|=k+1$ is called a $k$-dimensional simplex (or simply a $k$-simplex), where $k$ is a non-negative integer. Specifically, a vertex is a 00-simplex, an edge is a $1$-simplex, a triangle is a $2$-simplex, and so on.
For brevity, we use the term complex (or $d$-complex) instead of abstract simplicial complex throughout this article. The dimension of a complex $X$, denoted by $\dim(X)$, is the maximum dimension of all simplices in $X$, defined as
$\dim(X):=\max\{\dim(\sigma):\sigma\in X\},$
where $\dim(\sigma)=|\sigma|-1$. Note that if $\dim(X)=1$, $X$ can be considered as a graph whereas $\dim(X)=0$ represents a null graph.
For $0\leq j\leq\dim(X)$, the set of all $j$-simplices of $X$ is denoted by
$X_{j}:=\{\sigma\in X\mid\dim(\sigma)=j\}.$
For the remainder of the article, for $d\in\mathbb{N}$, the complex will take the form
$X:=\left(\bigcup_{k=0}^{(d-1)}X_{k}\right)\cup X^{d},$ (1)
where $\mathcal{X}_{0}:=\{1,2,\dots,n\}$, $X_{k}:=\{\{i_{0},\dots,i_{k}\}\mid 1\leq i_{0}<\dots<i_{k}\leq n\}$ for $1\leq k\leq d$, and $X^{d}\subseteq X_{d}$. Here, $X_{k}$ represents the set of all $k$-simplices on $\mathcal{X}_{0}$. In this model, $X$ is a complex with a complete $((d-1))$-dimensional skeleton.
Definition 2 (Weighted $d$-complex)
A weighted $d$-complex $(X,w)$ is a $d$-complex in which each member of $X^{d}$ is assigned a weight (or cost), typically represented as a real number. The weight of a $d$-simplex $\tau\in X^{d}$ is denoted by $w(\tau)$, where $w:X^{d}\to\mathbb{R}$ is a weight function.
We say that $\sigma,\sigma^{\prime}\in X_{(d-1)}$ are neighbors if $\sigma\cup\sigma^{\prime}\in X^{d}$, in which case we write $\sigma\sim\sigma^{\prime}$ (see Figure 2111Given a $d$-dimensional simplicial complex $X$, the figures illustrating $X$ display a selection of $(d-1)$-dimensional faces, rather than the exhaustive set of all $(d-1)$-simplices contained within $X$, in order to maintain visual concision and avoid clutter.). This concept is similar to the one introduced in [10].
[…]
Figure 2: Weighted 2-Complex and Neighbour
The degree of a simplex $\sigma\in X_{(d-1)}$ is denoted by
$\deg(\sigma)=\deg_{X}(\sigma):=\sum_{\tau\in X^{d}}\mathbf{1}_{\{\sigma\subset\tau\}},$
which counts the number of $d$-simplices containing $\sigma$. Note that a simplex $\tau\in X^{d}$ contributes to the sum if $\tau=\sigma\cup\{v\}$ for some $v\in\mathcal{X}_{0}\setminus\sigma$. The set of neighbors of $\sigma\in X_{(d-1)}$ is denoted by $S_{\sigma}$, that is,
$S_{\sigma}=\{\sigma^{\prime}\in X_{(d-1)}\mid\sigma^{\prime}\sim\sigma\}.$
$\{1,2,3\}$$\{2,3,4\}$$\{2,3,5\}$$d=2$$\deg(\{2,3\})=3$$S_{\sigma}=\{\{1,2\},\{1,3\},\{2,5\},\{3,5\},\{2,4\},\{3,4\}\}$$|S_{\sigma}|=d\deg(\sigma)$
Figure 3: Degree and Set of Neighbours
It follows that the number of elements in $S_{\sigma}$ is $d$ times $\deg(\sigma)$, i.e.,
$|S_{\sigma}|=d\deg(\sigma).$
Definition 3 ($d$-path and $d$-cycle)
A $d$-path in a $d$-complex $X$ is a sequence of distinct $(d-1)$-simplices $\sigma_{0},\sigma_{1},\dots,\sigma_{l}$ such that $(\sigma_{i}\cup\sigma_{i+1})\in X^{d}$ for all $0\leq i<l$ and the collection $\{\sigma_{i}\cup\sigma_{i+1}:0\leqslant i\leqslant l\}$ is a disjoint collection. The $d$-path is said to have length $l$ and is denoted by $P=(\sigma_{0},\sigma_{1},\dots,\sigma_{l})$. If the $d$-path starts and ends at the same $(d-1)$-simplex, it is called a $d$-cycle of length $l$.
1234567891011121314
[…]
In a weighted $d$-complex $(X,w)$, the shortest $d$-path between two vertices $\sigma_{1}$ and $\sigma_{2}$ is a $d$-path from $\sigma_{1}$ to $\sigma_{2}$ that minimizes the sum of the weights of the $d$-simplices along the $d$-path. The shortest $d$-path may not be unique. The weighted distance between two $(d-1)$-simplices $\sigma_{1},\sigma_{2}\in X_{(d-1)}$ is the sum of the weights of the $d$-simplices along the shortest $d$-path from $\sigma_{1}$ to $\sigma_{2}$. The weighted distance is denoted as $d_{w}(\sigma_{1},\sigma_{2})$.
If there is no $d$-path between $\sigma_{1}$ and $\sigma_{2}$, the weighted distance is defined to be infinity, i.e., $d_{w}(\sigma_{1},\sigma_{2})=\infty$.
2.2 Dijkstra’s Algorithm for Graphs
In this section, we briefly recall Dijkstra’s algorithm for graphs. Dijkstra’s algorithm [[11], Algorithm 2.3.5] finds the shortest path from a source $(d-1)$-simplex $\sigma_{0}$ (vertex) to all other vertices in a weighted graph $G=(V,E,w)$. The algorithm iteratively selects the $(d-1)$-simplex with the smallest known distance, relaxes its neighbours, and updates the shortest known distances. On a high level, the algorithm follows the following steps (1-3).
  1. 1.
    Initialize the distance to the source $d(s)=0$ and all other vertices $d(v)=\infty$.
  2. 2.
    Select the unvisited $(d-1)$-simplex $u$ with the smallest $d(u)$, and update the distances to its neighbors:
    $d(v)=\min(d(v),d(u)+w(u,v))$
  3. 3.
    Repeat until all vertices have been visited or the smallest distance is infinity.
3 Our Algorithm for Finding the Shortest Paths in a Weighted Simplicial Complex
In this section, we describe and analyse our novel algorithm. We also provide example of the workings of our algorithm.
Our algorithm is a greedy algorithm to find the shortest $d$-paths from a source $(d-1)$-simplex $\sigma_{0}$ to all other $(d-1)$-simplexes in a weighted complex $(X,w)$, where $w:X_{d}\to\mathbb{R}$ represents the weight function. Note that the weights have to be strictly positive for this algorithm to work. The algorithm works by iteratively selecting the $(d-1)$-simplexes with the smallest known distance and updating the distances to their neighbouring $(d-1)$-simplexes. The pseudo-code of our algorithm is given in Algorithm 1. But for ease of understanding, the high-level idea of our algorithm is described in the following steps.
Steps of Our Algorithm:
  1. 1.
    Input: A weighted $d$-complex with nonnegative weight function $w$ and a starting $(d-1)$-simplex $\sigma_{0}$. For any $x\in\mathcal{X}_{0}$, the weight of the $d$-simplex $\sigma_{0}\cup\{x\}$ is $w\left(\sigma_{0}\cup\{x\}\right)$; Let $w\left(\sigma_{0}\cup\{x\}\right)=\infty$ if $\sigma_{0}\cup\{x\}\notin X^{d}$.
  2. 2.
    Initialization:
    1. (a)
      Set the distance to the source $(d-1)$-simplex $\sigma_{0}$ as $t(\sigma_{0})=0$.
    2. (b)
      Set the distance to all other $(d-1)$-simplexes $\sigma\in X_{(d-1)}\setminus\{\sigma_{0}\}$ as $t(\sigma)=\infty$.
    3. (c)
      Mark all vertices as unvisited.
  3. 3.
    Iteration:
    1. (a)
      Select the unvisited $(d-1)$-simplex $\sigma_{1}$ with the smallest known distance $t(\sigma_{1})$. The choice of $\sigma_{1}$ may be more than one.
    2. (b)
      For each neighbor $\sigma_{2}$ of $\sigma_{1}$, if the $d$-path from $\sigma_{0}$ to $\sigma_{2}$ through $\sigma_{1}$ offers a shorter $d$-path than the current known distance, update the distance:
      $t(\sigma_{2})=\min(t(\sigma_{2}),t(\sigma_{1})+w(\sigma_{1},\sigma_{2}))$
  4. 4.
    Termination:
    1. (a)
      Mark $\sigma_{2}$ as visited. If all $(d-1)$-simplexes have been visited or the smallest known distance among the unvisited $(d-1)$-simplexes is infinity, the algorithm terminates.
    2. (b)
      The shortest $d$-path from $\sigma_{0}$ to each $(d-1)$-simplex is now known.
Input: A weighted $d$-complex with non-negative weight function $w$ and a starting $(d-1)$-simplex $\sigma_{0}$
Output: The shortest $d$-path from $\sigma_{0}$ to each $(d-1)$-simplex
1 Set the distance to the source $(d-1)$-simplex $\sigma_{0}$ as $t(\sigma_{0})=0$
2 Set the distance to all other $(d-1)$-simplexes $\sigma\in X_{(d-1)}\setminus\{\sigma_{0}\}$ as $t(\sigma)=\infty$
3 Mark all vertices as unvisited
4
5while there exists an unvisited $(d-1)$-simplex $\sigma_{1}$ with $t(\sigma_{1})<\infty$ do
6       Select an unvisited $\sigma_{1}$ minimizing $t(\sigma_{1})$
7       foreach unvisited neighbor $\sigma_{2}$ of $\sigma_{1}$ (i.e. $\sigma_{1}\cup\sigma_{2}\in X^{d}$) do
8             $t(\sigma_{2})\leftarrow\min\bigl{(}t(\sigma_{2}),\;t(\sigma_{1})+w(\sigma_{1}\cup\sigma_{2})\bigr{)}$
9
10      Mark $\sigma_{1}$ as visited
11
12return Distance function $t\colon X_{(d-1)}\to[0,\infty]\,$, where $t(\sigma)$ is the minimum total weight of a chain of $d$-simplices from $\sigma_{0}$ to $\sigma$.
13
Algorithm 1 Shortest $d$-paths in Simplicial Complexes
For an example of the application of this algorithm, see Figure 5.
[…]
Figure 5: Diagram explaining the execution of the algorithm using an example of finding $d_{w}(\{1,3\},\{4,7\})$ in the complex
3.1 Analysis of Algorithm 1
Correctness.
We begin by proving the correctness of our algorithm. For this, we state and prove the following theorem.
Theorem 3.1
Given a weighted $d$-complex with nonnegative weight function $w$ and a starting $(d-1)$-simplex $\sigma_{0}$, Algorithm 1 computes $d_{w}(\sigma_{0},\sigma_{1})$ for every $\sigma_{1}\in X_{(d-1)}.$
Proof
We prove that Algorithm 1 correctly computes the shortest $d$-path from the source $(d-1)$-simplex $\sigma_{0}$ to every other $(d-1)$-simplex in a weighted complex $(X,w)$, where $w$ is the weight function on the edges. We use Mathematical Induction to show that at each step of Dijkstra’s algorithm, the distance $t(\sigma_{1})$ is always the correct shortest $d$-path from the source $\sigma_{0}$ to $(d-1)$-simplex $\sigma_{1}$ for all vertices $\sigma_{1}\in X_{(d-1)}$.
At any step of the algorithm, let $\mathcal{S}$ be the set of $(d-1)$-simplexes whose shortest distance from the source $\sigma_{0}$ has been finalised. Initially, the set $\mathcal{S}=\{\sigma_{0}\}$, and the algorithm sets $t(\sigma_{0})=0$ and $t(\sigma_{2})=\infty$ for all $\sigma_{2}\neq\sigma_{0}$. Clearly, the distance from $\sigma_{0}$ to $\sigma_{0}$ is correct since $t(\sigma_{0})=0$. So, the algorithm correctly initialises the distances.
Assume that the algorithm has correctly computed the shortest $d$-path distances for all vertices in $\mathcal{S}$, where $|\mathcal{S}|=m$ for some $m\in\mathbb{N}$. Let $\sigma_{1}$ be the next $(d-1)$-simplex added to $\mathcal{S}$. By the algorithm’s rule, $\sigma_{1}$ is the unvisited $(d-1)$-simplex with the smallest tentative distance $t(\sigma_{1})$. We need to show that this choice of $\sigma_{1}$ guarantees that $t(\sigma_{1})$ is the correct shortest $d$-path from $\sigma_{0}$ to $\sigma_{1}$.
Any $d$-path $P$ from $\sigma_{0}$ to $\sigma_{1}$ must exit $\mathcal{S}$ before reaching $\sigma_{1}$. The induction hypothesis states that the length of the shortest $d$-path from $S$ to $\sigma_{1}$ is $t(\sigma_{1})$. The induction hypothesis and the choice of $\sigma_{1}$ also guarantee that a $d$-path visiting any $(d-1)$-simplex outside $\mathcal{S}$ and later reaching $\sigma_{1}$ has length at least $t(\sigma_{1})$. Hence $d_{w}(\sigma_{0},\sigma_{1})=t(\sigma_{1})$. Therefore, the algorithm correctly computes the shortest $d$-path to $\sigma_{1}$ and adds $\sigma_{1}$ to $\mathcal{S}$. The algorithm then proceeds to the next $(d-1)$-simplex in the same way, ensuring that at every step, the shortest $d$-paths from $\sigma_{0}$ to all vertices in $\mathcal{S}$ are correctly computed.
The algorithm terminates when $\mathcal{S}=X_{d-1}$, at which point $t(\sigma_{1})$ for each $(d-1)$-simplex $\sigma_{1}$ in $X_{d-1}$ represents the shortest $d$-path from $\sigma_{0}$ to $\sigma_{1}$.
Since the algorithm always selects the $(d-1)$-simplex with the smallest tentative distance at each step and updates distances optimally, the final distances $t(\sigma_{1})$ are indeed the correct shortest $d$-path distances from $\sigma_{0}$ to every $(d-1)$-simplex in $X_{d-1}$.
Hence, the correctness of the algorithm Algorithm 1 follows from Theorem 3.1.
Complexity.
Now we deduce the time and memory complexity of Algorithm 1. For clarity we denote the number of $d-1$-simplices by $n=|X_{d-1}|$ and the number of $d$-simplices by $m=|X_{d}|$. We will analyse the algorithm assuming that we use the binary heap data structure to store the distances. Now, if we use a binary heap structure for storing the current distances, then scanning the binary heap to find the smallest distance (extract-min) would take $\mathcal{O}(\log n)$ operations per extraction. Next, updating the entry (decrease-key) would take $\mathcal{O}(\log n)$ operations per iteration. Hence, the time-complexity of the algorithm is a sum of $n$ extractions and up to $m$ decrease-key operations. Therefore, the overall runtime of the algorithm is given by $\mathcal{O}((n+m)\log n)$.
For each $d-1$-simplex, we store the tentative distance from the starting point, which requires $\mathcal{O}(n)$ memory. Next, for each $d-1$-simplex, we need to store a boolean value that represents whether that simplex has been visited or not. This again requires $\mathcal{O}(n)$ memory. The algorithm also uses a priority queue to quickly extract an unvisited $d-1$-simplex with the smallest distance. The queue stores up to $n$ entries at once, each corresponding to a $d-1$-simplex and its associated distance. Therefore, the queue requires $\mathcal{O}(n)$ memory. Now, to find the neighbours of a given $d-1$-simplex, we must know all the $d$-simplices that contain the given $d-1$-simplex. It requires $\mathcal{O}(m)$ memory to store this adjacency information. This makes the total memory requirement of the algorithm to be $\mathcal{O}(m+n)$.
4 Conclusion
In this study, we have defined the concept of weight in a $d$-complex to facilitate our study of $d$-complex as an object in discrete mathematics and graph theory rather than a topological object. The concept of a $d$-path comes from this naturally. The main contribution of our paper is our novel algorithm to find the shortest path in a weighted $d$-complex. With its vast application in various fields, our algorithm is poised to play a crucial role in the study of complexes as mathematical objects with real-life applications. Our algorithm can be seen as the $d$-complex version of Dijkstra’s Algorithm for graphs. Analogy of a weighted $d$-complex with a weighted graph opens numerous questions in this field of study. The concepts introduced in this paper pave the path for a line of study in which simplicial complexes will be viewed as geometric computational objects rather than abstract entities.
This work extends beyond just providing a thorough analysis of our algorithm, this also indicates possible integration of simplicial complexes with various concepts of graph theory, specially weighted graphs.
One future aspect of study is the concept of a $d$-cycle (see Definition 3). The natural question to ask is, given $d$-complex how can we decide if the given $d$-complex contains a $d$-cycle or not. Now, if a $d$-cycle exists, how does one identify a $d$-cycle? The mono-dimensional version of these questions are of immense interest in the study of theoretical computer science and graph theory. The questions that we raised here are beyond the scope of this paper, hence we leave them as future directions.